# Experimental plasmonic sensing of malaria using an aluminum metasurface

### v1

@@ IV.2 Sensor calibration: bulk sensing
Figure 4: Sensor calibration using bulk organic solvents. (a) Measured transmission of the non-functionalized sensor in PBS ($n_{\text{eff}}=1.335$), ethanol ($n_{\text{eff}}=1.36$2), and isopropanol ($n_{\text{eff}}=1.377$). The inset shows a zoomed-in spline fitting of the measured data (solid lines) and the corresponding FEM simulated spectra (dotted lines). The bulk sensitivities are calculated for panel (b) with spectral sensing at resonance (including $n_{\text{eff}}=1.00$ for air) and panel (c) with intensity sensing off-resonance at $\lambda_{0}=660$ nm (the air resonance peak is too far away to be used). A sensitivity of 360(355)360\;(355) nm/RIU and 1.2053(1.2810)1.2053\;(1.2810) a.u./RIU are obtained based on the experimental data (FEM simulation) for spectral and intensity sensing, respectively.
@@
Bulk sensing measures the sensor’s response to changes in the bulk refractive index of the surrounding medium (e.g. buffer solutions or solvents with known refractive index $n_{\text{eff}}$). This stage is known as the bulk sensing calibration stage, and it forms a baseline for later pfLDH-specific measurements. We consider how the non-functionalized sensor responds to a bulk change in the refractive index $n_{\text{eff}}$. We injected two organic solvents into the wet chamber, ethanol ($n_{\text{eff}}=1.362$) and isopropanol ($n_{\text{eff}}=1.377$), one at a time. The sensor was completely rinsed with PBS and air-dried before injecting another liquid. The corresponding transmission profiles (for $\lambda=600$ nm to $725$ nm) are presented in Fig. 4 (a). We measured the spectral resonance shift of EOT peak (ii), $\delta\lambda_{R}=\lambda_{R}^{\text{eth/iso}}-\lambda_{R}^{\text{pbs}}$, and calculated the sensor sensitivity $S_{\lambda_{R}}$, as defined in Eq. (5). Experimental spectra are compared with the simulation spectra in the inset of Fig. 4 (a). Sensitivity calculations are presented in Fig. 4 (b). The sensor has a bulk sensitivity of $S_{\text{exp}}=360$ nm/RIU, which is close to the simulated sensitivity of $S_{\text{sim}}=355$ nm/RIU. These values are within the range of other aluminum [48, 49] and gold-based EOT sensors [63]. For intensity sensing, we fixed the wavelength at $\lambda_{0}=660$ nm and plotted the sensor transmittance against the solvent refractive index to obtain the intensity sensitivity, $S_{T}$. The results are presented in Fig. 4 (c). There is again a close match between the experimental ($S_{\text{exp}}=1.2053$ a.u./RIU) and simulation sensor sensitivity ($S_{\text{sim}}=1.2810$ a.u./RIU).
@@ IV.5 Control experiments
where $na_{\text{eff}}$ is the effective refractive index of the bound analyte layer, $nb_{\text{soln}}$ is the refractive index of the buffer, $da$ is the thickness of the analyte layer, $dc$ is the thickness of the biochemistry (3-GPS + Protein A + p-IgG) layer from the metal surface and $\delta_{d}$ is the probe depth of the localized SPs. For our sensor design, $S_{\lambda}=360$ nm/RIU and $S_{T}=1.2053$ a.u./RIU (see Fig. 4), $na_{\text{eff}}=1.45$, $nb_{\text{soln}}=1.335$, $da=8$ nm, $dc=19$ nm and $\delta_{d}$ ranges from 5010050-100 nm. Using these values in Eq. (9), the expected spectral shift, $\delta\lambda_{R}$, of the sensor is 4.25.34.2-5.3 nm and transmission shift, $\delta T$, is 0.0140.0180.014-0.018 a.u., equivalent to 0.140.180.14-0.18 a.u. when $T$ is normalized. These shifts, $\delta\lambda_{R}$ and $\delta T$, are consistent with those observed in our experiment at a high concentration of $250$ nM, as shown in Fig. 8(c) and (d), respectively.
@@ V Summary and Outlook
Among the different types of optical biosensors, plasmonic sensors have shown experimentally the possibility of detecting low concentrations of malaria biomarkers, such as pLDH spiked in buffer solution, or whole blood lysate. The best achieved LOD so far ranges from fM to pM based on a plasmon-enhanced fluorescence immunosensor. However, the achieved LOD based on fluorescence requires expensive equipment to detect the low light level of the fluorescence tags. For the SPR setup, the majority of demonstrated malaria plasmonic biosensors have been fabricated using gold and target pfHRP-2 as the malaria biomarker. The excellent plasmonic properties of gold – low optical losses in the visible and near-infrared regions and its high chemical stability are counteracted by its high cost, which limits large-scale commercialization. An attractive alternative to gold is aluminum. Aluminum is easy to fabricate and functionalize, and has properties that enable plasmon resonances in a broad optical band. Nano-engineered aluminum surfaces, such as the one studied in this work, have previously been proposed for use in biosensing applications [48, 49, 47].
We have shown that a low LOD is achievable using an aluminum metasurface as a plasmonic malaria sensor. The LOD of 1.3 nM, achieved using the Al-NHA sensor, is better than the previously reported LOD of 23.5 nM achieved using a gold-NHA [28]. Our results may encourage the design of cost-effective malaria sensors based on aluminum plasmonics for detecting the pLDH protein in real blood samples [83, 84]. Due to the similarity in amino acid sequences among plasmodium species, it has been shown that monoclonal antibodies that target pfLDH can also detect P. vivax and P. knowlesi [68]. Therefore, the sensor could potentially be used to detect any pLDH for biochemical research and point-of-care in the clinical field. Furthermore, the surface chemistry using an epoxy-silane linking protocol followed in this study is simple and effective. Both the Al-NHA and the surface chemistry developed are stable in liquid environments and can be used for a long time without degradation in sensor properties. In the future, we plan to test the sensor using intensity interrogation with a narrow bandwidth laser in the 650 - 670 nm range, to enhance the overall sensitivity and LOD to pM. Another direction to improve the LOD would be to replace the classical source of light with alternative quantum light sources [85] that could offer an improved LOD due to a decrease in shot noise [86].

### v2

@@ IV.2 Sensor calibration: bulk sensing
Figure 4: Sensor calibration using bulk organic solvents. (a) Measured transmission of the non-functionalized sensor in PBS ($n_{\text{eff}}=1.335$), ethanol ($n_{\text{eff}}=1.36$2), and isopropanol ($n_{\text{eff}}=1.377$). The inset shows a zoomed-in spline fitting of the measured data (solid lines) and the corresponding FEM simulated spectra (dotted lines). The bulk sensitivities are calculated for panel (b) with spectral sensing at resonance (including $n_{\text{eff}}=1.00$ for air) and panel (c) with intensity sensing off-resonance at $\lambda_{0}=660$ nm (the air resonance peak is too far away to be used). A sensitivity of 360(355)360\;(355) nm/RIU and 9.8876(9.2226)9.8876\;(9.2226) a.u./RIU are obtained based on the experimental data (FEM simulation) for spectral and intensity sensing, respectively.
@@
Bulk sensing measures the sensor’s response to changes in the bulk refractive index of the surrounding medium (e.g. buffer solutions or solvents with known refractive index $n_{\text{eff}}$). This stage is known as the bulk sensing calibration stage, and it forms a baseline for later pfLDH-specific measurements. We consider how the non-functionalized sensor responds to a bulk change in the refractive index $n_{\text{eff}}$. We injected two organic solvents into the wet chamber, ethanol ($n_{\text{eff}}=1.362$) and isopropanol ($n_{\text{eff}}=1.377$), one at a time. The sensor was completely rinsed with PBS and air-dried before injecting another liquid. The corresponding transmission profiles (for $\lambda=600$ nm to $725$ nm) are presented in Fig. 4 (a). We measured the spectral resonance shift of EOT peak (ii), $\delta\lambda_{R}=\lambda_{R}^{\text{eth/iso}}-\lambda_{R}^{\text{pbs}}$, and calculated the sensor sensitivity $S_{\lambda_{R}}$, as defined in Eq. (5). Experimental spectra are compared with the simulation spectra in the inset of Fig. 4 (a). Sensitivity calculations are presented in Fig. 4 (b). The sensor has a bulk sensitivity of $S_{\text{exp}}=360$ nm/RIU, which is close to the simulated sensitivity of $S_{\text{sim}}=355$ nm/RIU. These values are within the range of other aluminum [48, 49] and gold-based EOT sensors [63]. For intensity sensing, we fixed the wavelength at $\lambda_{0}=660$ nm and plotted the sensor transmittance drop, $\delta T_{\rm norm}$, against the solvent refractive index to obtain the intensity sensitivity, $S_{T}$. The results are presented in Fig. 4 (c). There is again a close match between the experimental ($S_{\text{exp}}=9.8876$ a.u./RIU) and simulation sensor sensitivity ($S_{\text{sim}}=9.2226$ a.u./RIU).
@@ IV.5 Control experiments
where $na_{\text{eff}}$ is the effective refractive index of the bound analyte layer, $nb_{\text{soln}}$ is the refractive index of the buffer, $da$ is the thickness of the analyte layer, $dc$ is the thickness of the biochemistry (3-GPS + Protein A + p-IgG) layer from the metal surface and $\delta_{d}$ is the probe depth of the localized SPs. For our sensor design, $S_{\lambda}=360$ nm/RIU and $S_{T}=9.8876$ a.u./RIU (see Fig. 4), $na_{\text{eff}}=1.45$, $nb_{\text{soln}}=1.335$, $da=8$ nm, $dc=19$ nm and $\delta_{d}$ ranges from 5010050-100 nm. Using these values in Eq. (9), the expected spectral shift, $\delta\lambda_{R}$, of the sensor is 4.25.34.2-5.3 nm and transmission shift, $\delta T$, is 0.140.180.14-0.18 a.u. These shifts, $\delta\lambda_{R}$ and $\delta T$, are consistent with those observed in our experiment at a high concentration of $250$ nM, as shown in Fig. 8(c) and (d), respectively.
@@ V Summary and Outlook
Among the different types of optical biosensors, plasmonic sensors have shown experimentally the possibility of detecting low concentrations of malaria biomarkers, such as pLDH spiked in buffer solution, or whole blood lysate. The best achieved LOD so far ranges from fM to pM based on a plasmon-enhanced fluorescence immunosensor. However, the achieved LOD based on fluorescence requires expensive equipment to detect the low light level of the fluorescence tags. For the SPR setup, the majority of demonstrated malaria plasmonic biosensors have been fabricated using gold and target pfHRP-2 as the malaria biomarker. The excellent plasmonic properties of gold – low optical losses in the visible and near-infrared regions and its high chemical stability are counteracted by its high cost, which limits large-scale commercialization. An attractive alternative to gold is aluminum. Aluminum is easy to fabricate and functionalize, and has properties that enable plasmon resonances in a broad optical band. Nano-engineered aluminum surfaces, such as the one studied in this work, have previously been proposed for use in other biosensing applications [48, 49, 47].
We have shown that a low LOD is achievable using an aluminum metasurface as a plasmonic malaria sensor. The LOD of 1.3 nM, achieved using the Al-NHA sensor, is better than the previously reported LOD of 23.5 nM achieved using a gold-NHA [28]. Our results may encourage the design of cost-effective malaria sensors based on aluminum plasmonics for detecting the pLDH protein in real blood samples [83, 84]. Due to the similarity in amino acid sequences among plasmodium species, it has been shown that monoclonal antibodies that target pfLDH can also detect P. vivax and P. knowlesi [68]. Therefore, the sensor could potentially be used to detect any species’ pLDH for biochemical research and point-of-care in the clinical field. Furthermore, the surface chemistry using an epoxy-silane linking protocol followed in this study is simple and effective. Both the Al-NHA and the surface chemistry developed are stable in liquid environments and can be used for a long time without degradation in sensor properties. In the future, we plan to test the sensor using intensity interrogation with a narrow bandwidth laser in the 650 - 670 nm range, to enhance the overall sensitivity and LOD. The sensor’s LOD may also be improved by using more specialised antibodies with lower $K_{D}$ values [68, 85]. Another direction to improve the LOD would be to replace the classical source of light with alternative quantum light sources [86] that could offer an improved LOD due to a decrease in shot noise [87].